show the equation used for the solution.
The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 34 in., find the dimensions of the rectangle.

Answer by anraymer(35)

L = 2 + 2W
P = 2L + 2W
P = 34 in.


34 = 2L + 2W
34 = 2(L + W)
17 = L + W
17 - W = L


L = 2 + 2W
17 - W = 2 + 2W
17 = 2 + 3W
15 = 3W
W = 5 in.


L = 2 + 2W
L = 2 + 2(5)
L = 2 + 10
L = 12 in.